Each of the "golden arches" at a McDonald's restaurant is in the shape of a parabola. Each arch is modeled by: h(x) = -x^2 + 6x, where h(x) is the height of the arch (in feet) at a distance x (in feet) from one side. Find the equation of the axis of symmetry, and how high is the arch at the axis of symmetry?
Answer by amarjeeth123(569)

Each arch is modeled by: h(x) = -x^2 + 6x, where h(x) is the height of the arch (in feet) at a distance x (in feet) from one side.
For axis of symmetry we have dh/dx=0
-2x+6=0
2x=6
x=3 is the axis of symmetry
Substituting x=3 we get h(3)=-9+18=9feet
The arch is 9 feet high at the axis of symmetry.